# Shared Spatial Memory Through Predictive Coding

### v1

@@ Methods
S1.4.2 Meta-controller architecture
@@
S1.4.3 Intrinsic curiosity module
@@
S1.4.5 Training configuration and optimization
@@
S1.5 Theoretical framework: Path integration and grid cell emergence
This section establishes a rigorous mathematical framework demonstrating how recurrent networks trained on predictive coding objectives naturally discover path integration dynamics and hexagonally-periodic spatial representations. We develop a principled theoretical argument showing that these solutions emerge from fundamental geometric constraints imposed by the task structure through three stages: (1) establishing preliminaries, (2) proving equivariance-driven path integration, and (3) deriving hexagonal symmetry from isotropy requirements.
S1.5.1 Preliminaries: Mathematical foundations
We formalize the mathematical structure underlying two-dimensional spatial navigation, establishing notation and dynamical equations that any successful path integration system typically incorporate.
@@
Definition 3 (Path integration dynamics).
@@
[…]
@@
S1.5.2 Equivariance under rigid body transformations
@@
S1.5.3 Emergence of sinusoidal phase encoding
@@
S1.5.4 Hexagonal symmetry from directional isotropy
@@
S1.6 Derivation of the variational bound for the predictive information bottleneck
This section establishes a rigorous mathematical framework for optimizing emergent communication protocols through variational inference. This framework is based on the well-known Information Bottleneck (IB) theory [1, 16], an elegant but intractable information theoretic solution, which is to optimize the inference quality under communication bandwidth constraints between two agents for fulfilling a specific mission. The resulting solution is the IB principle that balances the compression ratio for a certain level of inference quality and bandwidth consumption for such compressed information sharing. The core challenge lies in translating the IB principle into a computationally feasible training procedure. To use IB theory to solve our current problem at hand, its formulation naturally balances communication efficiency against predictive power by simultaneously minimizing the compression cost (rate) and maximizing the utility of transmitted messages for predicting the receiver’s future observations (distortion). However, direct optimization is fundamentally intractable, as it requires computing mutual information between high-dimensional neural representations governed by unknown probability distributions. We resolve this obstacle through variational inference, deriving tight upper for communication efficiency and lower bounds for prediction accuracy that transform the abstract IB objective into a fully differentiable loss function amenable to gradient-based optimization. This derivation proceeds as follows: we first establish notation and mathematical preliminaries, then derive tractable bounds for the rate term (Method S1.6.1), followed by bounds for the distortion term (Method S1.6.2), and finally synthesize these components into the unified variational information bottleneck (VIB) objective (Method S1.6.3).
@@
S1.6.1 Deriving a tractable upper bound for the communication rate
@@
S1.6.2 Deriving a tractable upper bound for the distortion
@@
where $h(p):=-\int p(x)\log p(x)\,dx$ denotes the Shannon entropy functional. Rearranging Ineq. (S56) gives $h(p(\cdot|z,S_{j}))\leq-\int p(O_{j}^{\prime}\mid z,S_{j})\log p_{\vartheta}(O_{j}^{\prime}\mid z,S_{j})\,dO_{j}^{\prime}$. Multiplying by 1-1 and taking expectations over $p(z,S_{j})$ on both sides, we have
@@
S1.6.3 Synthesis: Variational information bottleneck objective
@@
The VIB framework provides rigorous theoretical guarantees inherited from its variational foundation. Theorem 3 ensures that optimizing the tractable surrogate $\mathcal{L}_{\text{VIB}}$ drives down an upper bound on the true IB objective, guaranteeing improvement (under perfect optimization) toward the optimal rate-distortion trade-off. The tightness of this bound improves as the variational approximations become more accurate: when the encoder’s aggregate posterior $q(z)$ approaches the prior $p(z)$ and the decoder $p_{\vartheta}(O_{j}^{\prime}\mid z,S_{j})$ approaches the true conditional distribution, the inequalities in Lemmas 4 and 5 approach equality, yielding $\mathcal{L}_{\text{IB}}\to-H(O_{j}^{\prime}\mid S_{j})+\mathcal{L}_{\text{VIB}}$. Since the constant term $H(O_{j}^{\prime}\mid S_{j})$ is parameter-independent, optimizing $\mathcal{L}_{\text{VIB}}$ is equivalent to optimizing $\mathcal{L}_{\text{IB}}$ regardless of bound tightness.. Practically, this framework enables end-to-end learning of communication mechanisms through standard deep learning infrastructure—gradient-based optimization, mini-batch training, and GPU acceleration—without requiring explicit symbolic grounding, hand-crafted message spaces, or task-specific communication engineering. The learned protocols emerge purely from the objective of collaborative prediction under bandwidth constraints, embodying the social predictive coding principle: agents learn to exchange information that is maximally novel and decision-relevant from the receiver’s perspective, the essence of effective and efficient communication.
@@
Figure S4: Emergence of grid-like neural representations through self-supervised training. This figure contrasts the neural activation patterns of four representative units from the path integrator’s bottleneck layer before and after training, demonstrating the spontaneous formation of a grid-cell-like code. Before training, the units exhibit disorganized and non-periodic spatial firing fields ($F$), as shown in their spatial rate maps (top row) and corresponding spatial autocorrelograms (SACs, bottom row). The gridness scores are low (close to zero or negative), indicating a lack of hexagonal symmetry. After training on the self-motion prediction task, the same units develop highly structured, periodic firing patterns that tile the environment. The gridness scores are now significantly positive, and the SACs reveal a clear six-fold rotational symmetry, a defining characteristic of biological grid cells. This transformation illustrates how the predictive learning objective drives the self-organization of a stable, metric neural code for space.
@@
Figure S8: Gallery of emergent social place cells (SPCs). This figure shows representative units that selectively encode the partner’s location. Each row displays a single neuron’s activity across four conditions: (1) Self moving, peer static; (2) Peer moving, self static; (3) Both agents moving; (4) Both static. These SPCs exhibit strong, localized firing fields when the partner moves through a specific area (Condition 2), but remain largely silent in response to the agent’s own movement (Condition 1). This selective tuning to another’s position is the defining feature of social place cells.
@@
Figure S9: Gallery of emergent place cells (PCs). This figure shows representative units that function as classical place cells, selectively encoding the agent’s own location. The four conditions are shown for each neuron. These units display strong, stable firing fields when the agent itself traverses a specific location (Condition 1), but show negligible activation in response to the partner’s movement (Condition 2). This demonstrates a clear encoding of self-position, providing a stable allocentric representation for the agent.
@@
Figure S10: Gallery of special SPCs with mixed selectivity. This figure shows representative units, including “Relational Neurons”, that conjunctively encode both self and partner information. The four conditions are displayed for each unit. Unlike pure PCs or SPCs, these neurons show significant activation in response to both the agent’s own movement (Condition 1) and the partner’s movement (Condition 2). Firing is often maximal when both agents are moving (Condition 3), indicating that these cells encode a higher-order relational variable between the agents rather than a simple location.
@@
Figure S11: Robustness of the learned communication protocol to noise and scaling environmental complexity. This figure evaluates the HRL-ICM framework’s robustness to communication noise across a suite of procedurally generated mazes of increasing scale and complexity, shown in panels (a-d). Performance is compared against a “No Communication” baseline by measuring the average steps to completion (e) and the task success rate (f) under varying information loss rates (0-50%). The results demonstrate that HRL-ICM (blue) consistently solves tasks with higher efficiency and a significantly greater success rate. Crucially, its performance degrades gracefully as noise increases, whereas the baseline consistently fails. This robust advantage persists even in larger, more complex environments, highlighting the scalability of the learned protocol. Statistical significance is denoted as: * $p<0.05$, ** $p<0.01$, *** $p<0.001$.

### v2

@@ Methods
S2.4.2 Meta-controller architecture
@@
S2.4.3 Intrinsic curiosity module
@@
S2.4.5 Training configuration and optimization
@@
S2.5 Theoretical framework: Path integration and grid cell emergence
This section establishes a rigorous mathematical framework demonstrating how recurrent networks trained on predictive coding objectives naturally discover path integration dynamics and hexagonally-periodic spatial representations. We develop a principled theoretical argument showing that these solutions emerge from fundamental geometric constraints imposed by the task structure through three stages: (1) establishing preliminaries; (2) proving equivariance-driven path integration; and (3) deriving hexagonal symmetry from isotropy requirements.
S2.5.1 Preliminaries: Mathematical foundations
We formalize the mathematical structure underlying two-dimensional spatial navigation, establishing notation and dynamical equations that any successful path integration system typically incorporates.
@@
Definition 3 (Path integration dynamics[5]).
@@
$\mathbf{r}_{t+1}$ $=\mathbf{r}_{t}+\Delta\mathbf{r}_{t},\quad\mathrm{where}\quad\Delta\mathbf{r}_{t}=R(\theta_{t})\mathbf{v}_{t}\Delta t\penalty 10000\ .$ (S29)
@@
S2.5.2 Equivariance under rigid body transformations
@@
S2.5.3 Emergence of sinusoidal phase encoding
@@
S2.5.4 Hexagonal symmetry from directional isotropy
@@
S2.6 Derivation of the variational bound for the predictive information bottleneck
This section establishes a rigorous mathematical framework for optimizing emergent communication mechanism through variational inference. This framework is based on the well-known Information Bottleneck (IB) theory [1, 16], an elegant but intractable information-theoretic solution, optimizing inference quality under bandwidth constraints between two agents for fulfilling a specific mission. The resulting solution is the IB principle that balances the compression ratio for a certain level of inference quality and bandwidth consumption for such compressed information sharing. The core challenge lies in translating the IB principle into a computationally feasible training procedure. To use IB theory to solve our current problem at hand, its formulation naturally balances communication efficiency against predictive power by simultaneously minimizing the compression cost (rate) and maximizing the utility of transmitted messages for predicting the receiver’s future observations (distortion). However, direct optimization is fundamentally intractable, as it requires computing mutual information between high-dimensional neural representations governed by unknown probability distributions. We resolve this obstacle through variational inference, deriving tight upper bounds for communication efficiency and lower bounds for prediction accuracy that transform the abstract IB objective into a fully differentiable loss function amenable to gradient-based optimization. This derivation proceeds as follows: we first establish notation and mathematical preliminaries, then derive tractable bounds for the rate term (Method S2.6.1), followed by bounds for the distortion term (Method S2.6.2), and finally synthesize these components into the unified variational information bottleneck (VIB) objective (Method S2.6.3).
@@
S2.6.1 Deriving a tractable upper bound for the communication rate
@@
S2.6.2 Deriving a tractable upper bound for the distortion
@@
where $h(p):=-\int p(x)\log p(x)\,dx$ denotes the Shannon entropy functional. Rearranging inequality (S56) gives $h(p(\cdot|z,S_{j}))\leq-\int p(O_{j}^{\prime}\mid z,S_{j})\log p_{\vartheta}(O_{j}^{\prime}\mid z,S_{j})\,dO_{j}^{\prime}$. Multiplying by 1-1 and taking expectations over $p(z,S_{j})$ on both sides, we have
@@
S2.6.3 Synthesis: Variational information bottleneck objective
@@
The VIB framework provides rigorous theoretical guarantees inherited from its variational foundation. Theorem 3 ensures that optimizing the tractable surrogate $\mathcal{L}_{\text{VIB}}$ drives down an upper bound on the true IB objective, guaranteeing improvement (under perfect optimization) toward the optimal rate-distortion trade-off. The tightness of this bound improves as the variational approximations become more accurate: when the encoder’s aggregate posterior $q(z)$ approaches the prior $p(z)$ and the decoder $p_{\vartheta}(O_{j}^{\prime}\mid z,S_{j})$ approaches the true conditional distribution, the inequalities in Lemmas 4 and 5 approach equality, yielding $\mathcal{L}_{\text{IB}}\to-H(O_{j}^{\prime}\mid S_{j})+\mathcal{L}_{\text{VIB}}$. Since the constant term $H(O_{j}^{\prime}\mid S_{j})$ is parameter-independent, optimizing $\mathcal{L}_{\text{VIB}}$ is equivalent to optimizing $\mathcal{L}_{\text{IB}}$ regardless of bound tightness.. Practically, this framework enables end-to-end learning of communication mechanisms through standard deep learning infrastructure—gradient-based optimization, mini-batch training, and GPU acceleration—without requiring explicit symbolic grounding, hand-crafted message spaces, or task-specific communication engineering. The learned communication mechanisms emerge purely from the objective of collaborative prediction under bandwidth constraints, embodying the social predictive coding principle: agents learn to exchange information that is maximally novel and decision-relevant from the receiver’s perspective, the essence of effective and efficient communication.
@@
Figure S4: Emergence of grid-like neural representations through self-supervised training. Spatial activity patterns of nine representative artificial neurons from the path integrator’s bottleneck layer across training epochs (0, 5, 10, 20, 35). Top row of each pair shows firing fields ($F$); bottom row shows spatial autocorrelograms (SACs). Gridness scores ($G_{60}$, displayed below each SAC) increase from near-zero or negative values at initialization to highly positive values after training (up to 0.96), indicating emergence of hexagonal symmetry characteristic of biological grid cells. Unit numbers (#241, #63, etc.) indicate bottleneck layer indices. This progressive transformation demonstrates how predictive learning drives self-organization of structured spatial representations.
@@
Figure S8: Gallery of emergent social place cells (SPCs). Representative artificial neurons from the social processing network’s bottleneck layer that selectively encode the partner’s location. Each row displays a single artificial neuron’s spatial firing field across four experimental conditions: (1) Self moving, partner static; (2) Partner moving, self static; (3) Both agents moving; (4) Both static. Neuron numbers (e.g., #27, #34, #62, #80) indicate bottleneck layer indices. White lines show agent trajectories; red dots mark initial positions; heatmaps represent normalized firing rates (0-1 scale). Peak values indicate maximum activation for each condition. These SPCs exhibit strong, localized firing fields when the partner moves through a specific area (Condition 2), but remain largely silent in response to the agent’s own movement (Condition 1). This selective tuning to another’s position is the defining feature of social place cells.
@@
Figure S9: Gallery of emergent place cells (PCs). Representative artificial neurons from the bottleneck layer that function as classical place cells, selectively encoding the agent’s own location. Each row shows a single artificial neuron’s activity across the four experimental conditions (layout as in Fig. S4). Neuron numbers (e.g., #102, #118, #139, #208) indicate bottleneck layer indices. These units display strong, stable firing fields when the agent itself traverses a specific location (Condition 1), but show negligible activation in response to the partner’s movement (Condition 2). This demonstrates clear encoding of self-position, providing a stable allocentric representation for the agent.
@@
Figure S10: Gallery of special SPCs with mixed selectivity. Representative artificial neurons with conjunctive encoding of both self and partner information. Top row shows a bottleneck neuron (#145); remaining rows show relational head neurons (#3, #30, #35), illustrating that mixed selectivity emerges across multiple network layers. Each row displays activity across the four experimental conditions (layout as in Fig. S4). Unlike pure PCs or SPCs, these artificial neurons show significant activation in response to both the agent’s own movement (Condition 1) and the partner’s movement (Condition 2). Firing is often maximal when both agents are moving (Condition 3), indicating that these cells encode a higher-order relational variable between the agents rather than a simple location. The distinction between “Bottleneck” and “Relational” neurons reflects their position in different processing stages of the network architecture.
@@
Figure S11: Robustness of the learned communication mechanism to noise and scaling environmental complexity. This figure evaluates the HRL-ICM framework’s robustness to communication noise across a suite of procedurally generated mazes of increasing scale and complexity, shown in panels (a-d). Performance is compared against a “No Communication” baseline by measuring the average steps to completion (e) and the task success rate (f) under varying information loss rates (0-50%). The results demonstrate that HRL-ICM (blue) consistently solves tasks with higher efficiency and a significantly greater success rate. Crucially, its performance degrades gracefully as noise increases, whereas the baseline consistently fails. This robust advantage persists even in larger, more complex environments, highlighting the scalability of the learned communication mechanism. Statistical significance is denoted as: * $p<0.05$, ** $p<0.01$, *** $p<0.001$.
Appendix S4 Supplementary Videos
Video 1: Collaborative Mapping and Exploration. Here, we demonstrate the complete predictive coding framework for shared spatial memory in multi-agent maze navigation tasks. We compare baseline full broadcast communication against our HRL-ICM approach under limited communication budgets (20 rounds). The results show 44.7% faster task completion through adaptive communication strategies that avoid redundant information exchange, while baseline methods degrade after exhausting communication quotas early in exploration.
Video 2: Visualization of Emergent Social Place Cells. Here, we visualize the spontaneous emergence of social place cells that encode partner agent locations through predictive coding. We examine four scenarios with different movement patterns: self-movement with stationary peer, stationary self with moving peer, and synchronized movement. The analysis reveals distinct functional neuron types including pure place cells, pure social place cells, and special social place cells, combining both spatial references.
Video 3: Causal Intervention on Emergent Communication. Here, we demonstrate the functional significance of emergent communication through causal intervention experiments in collaborative search tasks. By injecting fake target information into the communication channel, we test whether communicated signals actually influence partner behavior. The intervention causes agents to navigate toward false targets, providing causal evidence that emergent communication mechanisms directly drive decision-making rather than being epiphenomenal.
Video 4: Full Framework - Collaborative Mapping and Exploration. Here, we present the complete demonstration of our predictive coding framework pipeline from egocentric vision to bird’s-eye-view map reconstruction. We show three increasingly complex tasks involving multi-agent coordination: single-agent BEV reconstruction, two-agent collaboration in multi-room environments, and three-agent coordination in large mazes. The framework achieves emergent division of labor through bandwidth-efficient communication (4-128 bits/step).
Video 5: Spontaneous Formation of Grid-Cell-like Representations. Here, we visualize the developmental trajectory of hexagonal grid cell representations emerging through self-supervised motion prediction without explicit supervision. The video shows evolution from random neural activity (Epoch 0) to organized hexagonal tessellations with 60-degree rotational symmetry (Epochs 20-150). Grid cell emergence correlates with dramatically improved trajectory prediction accuracy, demonstrating that predictive uncertainty minimization principles sufficiently explain mammalian entorhinal cortex spatial coding.